# WikiHint: A Human-Annotated Dataset for Hint Ranking and Generation

### v2

@@ Abstract.
The use of Large Language Models (LLMs) has increased significantly with users frequently asking questions to chatbots. In the time when information is readily accessible, it is crucial to stimulate and preserve human cognitive abilities and maintain strong reasoning skills. This paper addresses such challenges by promoting the use of hints as an alternative or a supplement to direct answers. We first introduce a manually constructed hint dataset, WikiHint, which is based on Wikipedia and includes 5,000 hints created for 1,000 questions. We then finetune open-source LLMs such as LLaMA-3.1 for hint generation in answer-aware and answer-agnostic contexts. We assess the effectiveness of the hints with human participants who answer questions with and without the aid of hints. Additionally, we introduce a lightweight evaluation method, HintRank, to evaluate and rank hints in both answer-aware and answer-agnostic settings. Our findings show that (a) the dataset helps generate more effective hints, (b) including answer information along with questions generally improves quality of generated hints, and (c) encoder-based models perform better than decoder-based models in hint ranking.
Hint Dataset, Hint Evaluation, Hint Generation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; July 13–18, 2025; Padova, ITisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and rankingccs: Information systems Evaluation of retrieval results
@@ 1. Introduction
Figure 1. Hints for a question from the WikiHint dataset, with their corresponding rankings (1 being the highest and 5 the lowest) which let users find the answer. Note that the opposite arrangement of the hints (from 1 to 5) would make the answer finding task easiest.
@@
In this paper, we propose the first manually curated dataset for the Automatic Hint Generation (HG) task, called WikiHint  [^2] https://github.com/DataScienceUIBK/WikiHint, which has been constructed and designed for hint generation, ranking and evaluation. Figure 1 shows a question from the WikiHint dataset, with its hints and the corresponding ranks reflecting the usefulness of each hint. We explore the performance of various LLMs in generating hints across different scenarios, including vanilla and finetuned models. We also examine the quality of the generated hints using both answer-aware and answer-agnostic approaches. Finally, we assess the effectiveness of a novel evaluation method for hint ranking, called HintRank, and compare it with other automatic evaluation techniques. Ranking hints allows arranging them either from easiest to hardest which can be used for selecting the most useful hints for a given question, or from hardest to easiest (as shown in Fig. 1). The latter arrangement could find applications in quizzes, knowledge assessment scenarios, or in certain entertainment use cases (e.g., in games).
@@ 3.1.1. Question Selection
readability The readability score.
familiarity The familiarity score.
@@ 3.2.1. Amazon MTurk
familiarity The familiarity score.
difficulty The difficulty level of the answer.
@@ 4. Evaluation Approaches
In this section, we first discuss the evaluation metrics we use to assess the quality of hints, followed by a detailed introduction to HintRank, our proposed automatic ranking method.
@@ 5.3. Model Performance
To evaluate the hint generation capabilities of LLMs across different scenarios, we examine four approaches: Vanilla-wA, Vanilla-woA, FTwA, and FTwoA, where FT stands for Finetuned, wA denotes With Answer, and woA means Without Answer. We test these models on the WikiHint test set to analyze the impact of finetuning, the inclusion of answers in the prompt, and their overall effectiveness in generating informative hints.
@@ 6. Conclusions
In the future, we plan to generate personalized hints tailored to the knowledge of askers. The main challenge here will be to develop appropriate datasets and solutions for user profiling.

### v3

@@ Abstract.
The use of Large Language Models (LLMs) has increased significantly with users frequently asking questions to chatbots. In the time when information is readily accessible, it is crucial to stimulate and preserve human cognitive abilities and maintain strong reasoning skills. This paper addresses such challenges by promoting the use of hints as an alternative or a supplement to direct answers. We first introduce a manually constructed hint dataset, WikiHint, which is based on Wikipedia and includes 5,000 hints created for 1,000 questions. We then finetune open-source LLMs for hint generation in answer-aware and answer-agnostic contexts. We assess the effectiveness of the hints with human participants who answer questions with and without the aid of hints. Additionally, we introduce a lightweight evaluation method, HintRank, to evaluate and rank hints in both answer-aware and answer-agnostic settings. Our findings show that (a) the dataset helps generate more effective hints, (b) including answer information along with questions generally improves the quality of generated hints, and (c) encoder-based models perform better than decoder-based models in hint ranking.
Hint Dataset, Hint Evaluation, Hint Generation, Hint Ranking
journalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730284isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Retrieval models and rankingccs: Information systems Evaluation of retrieval results
@@ 1. Introduction
Figure 1. Hints for a question from the WikiHint dataset, with their corresponding rankings (1 being the highest and 5 the lowest) which let users find the answer. Note that the opposite arrangement of the hints would make the answer finding task easiest when hints are read from rank 1 to 5.
[…]
@@
In this paper, we propose the first manually curated dataset for the Automatic Hint Generation (HG) task, called WikiHint  [^2] https://github.com/DataScienceUIBK/WikiHint, which has been constructed and designed for hint generation, ranking, and evaluation. Figure 1 shows a question from the WikiHint dataset, with its hints and the corresponding ranks reflecting the usefulness of each hint. We explore the performance of various LLMs in generating hints across different scenarios, including vanilla and finetuned models. We also examine the quality of the generated hints using both answer-aware and answer-agnostic approaches. Finally, we assess the effectiveness of a novel evaluation method for hint ranking, called HintRank, and compare it with other automatic evaluation techniques. Ranking hints allows arranging them either from easiest to hardest which can be used for selecting the most useful hints for a given question, or from hardest to easiest (as shown in Fig. 1). The latter arrangement could find applications in quizzes, knowledge assessment scenarios, or in certain entertainment use cases (e.g., in games).
@@ 3.1.1. Question Selection
readability The readability score of the question.
familiarity The familiarity score of the question.
@@ 3.2.1. Amazon MTurk
familiarity The familiarity score of the answer.
[…]
@@ 4. Evaluation Approaches
In this section, we begin by introducing the evaluation metrics used to assess the quality of hints. These metrics are available through the HintEval (Mozafari et al., 2025) framework, which provides a standardized setup for hint evaluation [^6] https://github.com/DataScienceUIBK/HintEval. Following this, we present HintRank, our proposed method for automatic hint ranking.
@@ 5.3. Model Performance
To evaluate the hint generation capabilities of LLMs across different scenarios, we examine four approaches including Vanilla-wA, Vanilla-woA, FTwA, and FTwoA, where FT stands for Finetuned, wA denotes With Answer, and woA means Without Answer. We test these models on the WikiHint test set to analyze the impact of finetuning, the inclusion of answers in the prompt, and their overall effectiveness in generating informative hints.
@@ 6. Conclusion
In the future, we plan to generate personalized hints tailored to the knowledge of users. The main challenge here will be to develop appropriate datasets and solutions for user profiling.